Grade 12General Physics

The time period of a mass suspended from a spring is T. If the spring is cut into 4 equal parts and the same mass is suspended from one of the parts then the new time period will be
(A) T/2
(B) T
(C) T/4
(D) 2T
Please help........................

MANAV S
8 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
By cutting spring in 4 equal parts
K' = 4k
Now
Time period = 2\pi \sqrt (m/k)
Now
New time period will be T' = T/2
Regards
Arun (askIITians forum expert)